OpenStudy (anonymous):

which polynomial function has the most roots? A) f(x)=x^12-1 B) g(m)=158m^2 C) f(a)=15z^4+a^3-3a^2 D )g(n)=2n^3+3n^6

OpenStudy (zehanz):

Do you know how to find roots?

OpenStudy (anonymous):

no

OpenStudy (zehanz):

You have to solve the equation f(x)=0 (or g(m)=0 etc.) So to see how many roots f has, you need to solve x^12-1=0

OpenStudy (zehanz):

I will help you with this one: x^12 - 1 = 0 means: x^12 = 1 (because 1 - 1 = 0). Now there is at least one number that, when raised to the power 12, equals 1. What number is that?

OpenStudy (anonymous):

im not quit sure

OpenStudy (zehanz):

It is 1. See: 1*1*1*...*1 = 1. Even if you multiply 1 twelve times with itself, it always remains 1! So 1 is a root of this equation. But there is another root! This means there is another number, that when raised to the 12th power, equals 1. (it is also a very "easy" number...)

OpenStudy (anonymous):

okay now im starting to understand